OpenStudy (lukecrayonz):

Can someone explain how to find the period of a trig function? Example problem: y=2cos(x/3)

OpenStudy (turingtest):

For trig equations of the form\[A\cos(Bx+C)\]the period T is given by\[T=\frac{2\pi}{|B|}\]identify your B and plug it into the formula for the period.

OpenStudy (lukecrayonz):

So your b would be what for that problem?

OpenStudy (turingtest):

I'm not going to tell you straight away. Ask yourself: What is A in this formula? What is C? When you identify those, the remaining number will be B. It should be obvious when you figure it out.

OpenStudy (turingtest):

want a hint?

OpenStudy (lukecrayonz):

Yes D:

OpenStudy (lukecrayonz):

A is 2 haha, i have no idea since its a fraction :S

OpenStudy (turingtest):

\[A\cos(Bx+C)\]Yours can be written as \[2\cos(\frac{1}{3}x+0)\]Does that make it more clear?

OpenStudy (lukecrayonz):

Yes :P Now i feel stupid.

OpenStudy (turingtest):

Please don't. New concepts are like that. Happens to me almost every day ;)